A particular medication dosage is 45.0 mg/kg of body weight. What is the mass in mg of the medication a child weighing 33.9 lb. should receive?



1
Expert's answer
2021-08-28T06:27:08-0400

At first, pounds should be turned into kilogrammes:


m=33.9*0.45359237=15.376781343(kg)~15.38(kg)


Dosage=45.0*15.38=692.1(mg)